A ball was kicked at an angle of 15 degree above the horizontal. The speed is 30 m/s at horizontal dimension (initial). Neglecting air resistance, the projectile will travel at what horizontal distance before striking the ground? Hint: Find first the time before its strikes the ground. Bonus: Get the time as it reaches its peak, the maximum height.


1
Expert's answer
2021-12-20T10:29:53-0500

Given:

"\theta=15^\circ"

"v_0=30\:\rm m/s"

"g=9.8\:\rm m/s^2"


The range of projectile

"R=\frac{v_0^2\sin2\theta}{g}=\rm\frac{(30\:m/s)^2\sin30^\circ}{9.8\: m/s^2}=46\: m"

The time of movement to the highest point of trajectory

"t=\frac{v_0\sin\theta}{g}=\rm\frac{30\: m/s*\sin15^\circ}{9.8\: m/s^2}=0.79\: s"